the length of a a rectangular garden is two less than three times the width. Find the length if the perimeter of the garden is 36 feet.
Answer by tara0066(31)

the length of a a rectangular garden is two less than three times the width. Find the length if the perimeter of the garden is 36 feet.
Perimeter: 2(Length)+2(Width)
Length:3x-2
Width:x
2(3x-2)+2(x)=36 Distribute the 2
(2)(3x)-(2)(2)+2x=36
6x-4+2x=36 Simplify
8x-4=36 Add 4 to both sides
8x-4(+4)=36(+4)
8x=40 Divide both sides by 8
x=5
Length: 3(5)-2=15-2=13
Width: 5